OpenStudy (anonymous):

x^2-2x=2...i have x^2-2x(1)^2=2(1)^2=x(x-2)^2=2 but it doesnt look rite

OpenStudy (anonymous):

^^ completing the squares

OpenStudy (anonymous):

\[x^2-2x=2\]\[x^2-2x+1=3\]\[(x-1)^2=3\]make sense?

OpenStudy (anonymous):

not really...so when you take the half of the number and square it then you dont use that #??

OpenStudy (anonymous):

you know when i did it i still had the 2 and did you already add the 1 to the 3??

OpenStudy (tyteen4a03):

@Blaque23 I showed you how to complete the square in another question you asked.

OpenStudy (anonymous):

i tried that with this problem and i put the answer i had it wasnt the same as what he just gave me...where did i go wrong??

OpenStudy (tyteen4a03):

By completing the square for x^2-2x, the correct answer is (x - 1)^2 - 1.

OpenStudy (anonymous):

so ur pretty much factoring it out and instead of doing that i combined them??

OpenStudy (anonymous):

ok so i have x^2+4x+1=0...x^2+4x=-1=x^2+4x+2=-1(2)=(x+2)^2=-2=-2+i√2...-2-i√2...did i do it rite this time??